2020 August/September Restart OUTLINE

August 24-September 4, 2020 - Engineering On-Sky Restart Block

General Scope of Activities

  • Closed Dome/Daytime Instrument Checks (each instrument must be released before night-time operations)
  • Telescope Wake-Up
  • Telescope Infrastructure
    • AGW1
    • SX M3
  • Checkout of Facility Instruments
  • LBTI Restart (1 x 0.5 nights)
  • AO (SX) Restart (2 x 0.5 nights)

Key Dates & Information

  • August 14 - DONE restart ICS
  • August 17/18 - DONE LUCI-1 cold
  • August 17 - DONE - start cooling LUCI-2 on telescope (should be available by Aug 25 - DAYTIME)
  • August 18 - DONE LUCI-1 cold check in lab before going on telescope - DAYTIME
  • August 19 - DONE Install LBCs on telescope and start pumping - DAYTIME
  • August 20 - DONE M3 SX installed (was Aug 21)
  • August 21 - DONE MODS 1/2 cool-down - DAYTIME
  • August 20-21 - Functional checkouts of LBC - DAYTIME
    • LBC-B Science Controller issue - NOT RELEASED
  • August 21 - DONE Functional Checkout of MODS
  • August 21 - DONE SX AdSec cover removed
  • August 22 - DONE MODS mask exchange for Restart
  • August 24 - DONE Projected date of LUCI-2 cool enough for functional testing
  • August 25 - DONE LUCI-2 passed functional check
  • August 25 - DONE Projected date for AGW-1 Install
  • August 27 - DONE LUCI-1 install - DAYTIME
  • August 27 or 28 - DONE LUCI-1 functional checkout (can be released for science after)
  • August 29-31 - LUCI flexure, wavelength calibration checks
  • Sep 01 - DONE target date for release of SX AO for use on-sky
  • Sep 01 - DONE first OSCO
  • August 30 - DONE LUCI-1 Mask Insertion, LUCI-2 extracton (AUX 2) - DAYTIME
  • August 31 - DONE LUCI-2 cabinet insertion (AUX 2) - DAYTIME
  • August 31 - DONE Target for full release of AO for science - May be delayed to Sep 01 - SOUL/SX released for on-sky work
    • 2 x 0.5 nights separated on-sky
  • TBD - DX AdSec Install

Timetable Update:

  • Sep 1 - complete LUCI-1 RotZP /AGW transform, LUCI dither tests, continue through facility insturment checks.
  • Sep 2 - closed - no backup power available
  • Sep 3
    • 1st half LBTi on-sky
    • 2nd half AO/SOUL/LUCI basic checkout - ONLY if LUCI-1 critical priority work completed
  • Sep 3
    • Contingency for LBTI/AO (should Sep 2nd be wiped out)
    • 1st half AO 2nd part of on-sky checkout
    • 2nd half - continue with facility instrument tests (LBC superfocus, mixed mode LBC-B/MODS-2)
  • Sep 7 - INAF request for ESM (AO/SX)

Personnel

  • Night Activities
    • Aug 21 - Aug 24 - Barry, John H.
      • Telescope Wakeup
      • Start of LBC wakeup
      • Start of MODS wakeup
        • Rotator ZP and Transform verification Aug 23 or Aug 24
    • Aug 25 - Aug 29 - Olga, John H. (as needed)
      • Finish LBC wakeup
      • Finish MODS wakeup
      • SX M3 alignment
      • AGW-1 alignment and checks
      • PEPSI FPU alignment
      • LUCI-2 basic on-sky checks
      • LUCI-1 basic on-sky checks (if released)
    • Aug 30 - Sep 2 - Jenny, John H. (as needed)
      • LUCI-1 checks continue- AGW-1 work
      • AGW1/2 2nd set of transforms
      • Confirmatory AGW transform collection both LUCIs (Sep 2)
      • LUCI-1/SX-AO SOUL (0.5 night) - ready for all AO activities (Aug 31)
      • AO Sep 2 - D. Miller
    • Sep 3 - Sep 4 - Jenny, Barry (AO scripting work), John H (as needed)
      • LUCI-1/SX-AO SOUL (0.5 night) - highest priority if AO not completed earlier
      • Complete tasks not completed
      • LBTI (0.5 night) Sep 3 (Sep 4 preferred)
      • AO Sep 3 - D. Miller
      • AO Sep 4. - G. Taylor
    • Sep 5 - D-night - Jenny/Barry
      • Rollover for any remaining restart activities
      • AO highest priority
      • AO G. Taylor
  • Daytime Activities
    • Aug 19-22 - LBC/MODS functional checks - Olga
    • Aug 22-28 - LUCI functional checks - Dave T./Jenny/Barry

Closed Dome/Daytime Work

  • LBC (2 hrs)
    • DONE Filter Functional checks
    • DONE Rotator Functional checks
    • DONE Test Bias Exposures
  • MODS
  • LUCIs (5-6 days total)
    • Basic Checks needed for release for engineering (and AO) work - 1 day i n total
      • DONE non-MOS Functional Checkout on telescope L1 (cold) - 0.5 day
      • DONE non-MOS Functional Checkout on telescope L2 (cold) - 0.5 day
      • DONE MOS Functional Checkout on telescope L1 (cold) - 0.5 day
      • DONE MOS Functional Checkout on telescope L2 (cold) - 0.5 day
      • DONE PFS alignment and Camera refocus L1 - 1hr
      • DONE PFS alignment and Camera refocus L2- 1hr
    • Grating Wavelength Calibration - 3-5hr take data, 1-2 days process data
      • G210 L1
      • G210 L2
      • G200 L1
      • G200 L2
      • G150 L1 (low priority)
      • G040 L2 (lowest priority)
    • Flexure Checks - 3 days
      • G210/N1.8 - L1/L2
      • G210/N3.75 - L1/L2
      • G200/N1.8 - L1/L2
      • G150/N1.8 - L1
      • N1.8 imaging - L1/L2
      • N3.75 imaging - L1/L2
      • N30 Imaging - L1/L2
  • LBTI (2 days)
    • DONE Internal checkout (0.5 day)
    • DONE AdSec check out (1 day) - currently scheduled for Sep 2 (could be done Sep 1)
  • AO (3-4 days)
    • DONE 4D alignment (2 x 0.5 days or 1 day)
    • DONE AO WFS checkout (internal)
    • DONE AO flats - (2 x 0.5 sharaed with LUCI checkout)
    • DONE Full AO checkout - (2 x 0.5 days - shared with LUCI checkout)
    • DONE Full AO Checkout with LUCI - including NCPA (2 x 0.5 day - shared with LUCI Cabex's)
    • DONE ARGOS cal unit checkout
  • PEPSI
    • DONE Internal Operations (on-sky with solar telescope already)
    • Test presets with MODS/PEPSI mixed mode

List of On-Sky Tasks

  • DONE Telescope Wake-up (1 hr) - Aug 21
  • Telecope Infrastructure
    • SX M3 Alignment - Critical (3 hrs) - requires a LUCI or LBTI or PEPSI
      • DONE Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout
      • DONE Pupil wobble (2 hrs) - requires LUCI-1/AGW-1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
      • Field Aberration (2 hrs) - requires AGW-1 to guide, collimate, SH spots well illuminated, hotspot centered, ROT ZP found - can be done before AGW Rot ZP and Transform
    • DX AdSec Install/On-sky verification (2hrs - TBD)
  • LBC (high) - LBC-Red shims replaced Sep 04 - needed to redo some tasks after
    • Collimate WRS/DOFPIA - high (4 x 0.25 hrs)
    • Superfocus sequence -PA=0,180 high (1 hr)
    • DONE (Both) Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr)
    • Check AdSec SX issues with LBC-Blue in the beam - there may be an issue with elevations of 30 (G. Brusa - daytime and night checks)
    • Test Science OB - guiding - medium (1 hr)
      • DONE Sidereal
      • Non-Sidereal OB
        • DONE (Red Only) non-dither
        • DONE (Red Only) dithered
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations
      • LBC-R data collected UT 26 Aug
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete)
      • as of 2020-Sep-1, we have LBCR flats @ r-SLOAN, R-BESSEL, i-SLOAN, Y-FAN and F972N20
      • We need LBCR: z-SLOAN, I-BESSEL and V-BESSEL and all LBCB filters
    • DONE (Red Only) on-sky rotational trailing tests
    • TMS with LBC/DX (Heejoo and John H.)
    • Test Active Optics issues (TBD)
  • MODS (high)
    • check MODS1/MODS2 guide probe wrt/rotator center (low priority - MODS were not removed nor AGWs removed)
      • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
      • Based on on-sky observing the hotspot and SH spots look to be OK - MODS was not removed or worked on in any way that should affect the AGWs
    • DONE MODS Rotator Center measure (can use same field as for astrometric masks - StoneO or StoneA) - requires pointing mask - high (0.5hr)
    • DONE AGW 5/6 - Measure Rotator zero point (mask alignment, masks for StoneO and StoneA) - high (0.5 hr) - mask exchange required (Aug 23)
      • MOS MMS files:
        • Field RA (J2000) Dec (J2000) MASK ID
          StoneO 21 37 22.232 00 31 54.19 535231
          StoneA 00 56 20.127 -01 14 12.57 571722
      • M1: dPA = -0.157 deg
      • M2: dPA = -0.011 deg
    • Not Needed AGW Transform data - ~1" seeing - high (2hours to collect data on-sky, time during day to analyze, ~0.5 hr-1hr to verify on-sky) - low priority
      • Down the road it might be useful to improve the M1 rotator angle but this is not a high priority
      • AGW transform data to be taken ONLY if run out of other tasks (could be used as training for analyzing WFS data)
    • Spectrophotometric standard all modes - medium
      • DONE Grating
      • DONE Prism
    • non-sidereal guiding
      • DONE Imaging
      • Long-slit - low priority
    • Test Science OB
      • DONE longslit
      • imaging
        • DONE Test without syncoffset and large dither (+/-45")
          • MODS2B slowed down and the offset occurred before exposure completed
        • DONE Test with synoffset and large dither (+/-45")
          • sides got out of xync and timed out - should have not timed out
    • DONE Dithering Repeatability with Pointing Mask to check DET XY
    • Skyflats
    • DONE Blind offset tests
  • LUCIs (high)
    • Rotator center - L1/L2
      • DONE L1
      • DONE L2
    • AGW1 Alignment (LUCI-1) - install planned Aug 22 (1 night) Overview of AGW procedure
      • DONE verify SH spots illuminated (0.5 hr)
      • DONE verify hotspot center (1 hr)
      • DONE rotator zeropoint (requires MOS mask - Aug 30th) - (0.5 hr)
        • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
      • DONE Longslit Method
        • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
      • DONE MOS LMS Files
Field RA (J2000) Dec (J2000) MASK ID
StoneO (old) 21 37 27.784 00 31 50.91 930819
StoneA (old) 00 56 16.168 -01 14 34.96 974550
StoneO (new) 21 37 31.791 00 31 55.010 903033
StoneP (new) 22 35 49.616 01 31 33.950 933322
StoneA (new) 00 54 02.050 -00 32 34.460 921934
    • AGW2
      • DONE rotator zeropoint (requires MOS mask - Sep 2)
      • DONE Longslit methtod for Rotator ZP (~0.12 degrees)
        • verify transform data - taken in parallel with LUCI-1/AGW1 verification - Sep 2nd or later - NOT NEEDED
    • DONE (G200 zJHK) Longslit check (dither along slit, offsets and dithering image tests) - high (1 hr)

    • Imaging check
      • DONE dither in field (0.5hr) -
      • DONE dither out of field (0.5hr) - M52 K and some J-band
    • non-sidereal Imaging Check
      • DONE dither in field (0.5hr) - Comet 29/P ~30 minutes K,H,J
      • DONE / N dither out of field (0.5hr) - worked LUCI-2 only, errors accumulated over time in Bino mode IT 8179
    • non-sidereal check longslit - low (0.5 hr)
      • Science-like OB for G210
      • Science-like OB for G200 (combined with NS longslit below)
    • Twilight flats
  • Mixed Mode Tests
    • LBC-Blue/LUCI-2 NS Mixed Mode verification (0.5hr) - awaiting repairs to LBC-B, or LUCI-1 installed
    • LBC-Red/MODS-1 (0.5-1hr)
  • LBTI (0.5 night)
    • Close AO loop on bright star.
    • Perform focus sequence in H band, confirm in L band.
    • Check for remaining NCPA in H band, L band, confirm PSF is good.
    • Test observing script for standard LMIRCam sequence.
    • Take OSCO-like data set.
  • AO (2 x 0.5 night - need not be consecutve)
    • System checkout (0.5 night - any half)
      • Diffraction limited verification
        • DONE Verify Bin 1
        • DONE Verify Bin 4 - some elongation, hit R=18
        • Test Science scripts (5 hrs)
          • DONE sideereal
          • Non-sidereal
        • DONE OSCO (0.5hr)
      • ESM verification
        • DONE Test science scripts (2 hr)
    • DONE Science check (0.5 night) - 1st half
  • PEPSI (0.5 night - low priority)
    • Verify PFU alignment

Instrument Released for Partner Science

  • DONE MODS
    • Caveats
      • Some refinement of focus may be needed
      • IT 6386 - sequecner issue of MODS2-Blue - intermittment - not clear if frequency is increasing, although it happens at the start of every night
      • SYNCOFFSET for image dithering needs work - if one MODS or MODS channel is slow, the SYNC offset will hang, removing SYNCOFFSET may result in telescope moving before an exposure finishes if one channel is too slow, as well as violating co-pointing limits (known issue)
      • MODS1B as of 08/27/20 is offline - repair of microlynx controller needed for grating turret - IT 8171
  • DONE LUCIs
    • Caveats
      • New Flexure and Wavelength tables not updated (flexure should not have changed with LUCI-2)
      • Co-pointing limits have to be diasbled for big dithers (PCS issue)
  • LBTI

Rainy Night Tasks

  • Pending LUCI closed dome tasks
  • Pending AO closed dome tasks
  • Pending LBTI closed dome tasks
  • TMS testing

-- BarryRothberg - 22 Jul 2020
Topic attachments
I Attachment Action SizeSorted ascending Date Who Comment
luci1.29.stnA2020.lmslms luci1.29.stnA2020.lms manage 6 K 18 Aug 2020 - 19:44 BarryRothberg  
luci1.32.stnP2020.lmslms luci1.32.stnP2020.lms manage 6 K 18 Aug 2020 - 19:43 BarryRothberg  
luci2.18.StoneA.lmslms luci2.18.StoneA.lms manage 6 K 18 Aug 2020 - 19:41 BarryRothberg  
luci1.27.stnO2020.lmslms luci1.27.stnO2020.lms manage 8 K 18 Aug 2020 - 19:42 BarryRothberg  
luci2.15.StoneO.lmslms luci2.15.StoneO.lms manage 8 K 18 Aug 2020 - 19:39 BarryRothberg  
mods.37.StoneA.mmsmms mods.37.StoneA.mms manage 8 K 18 Aug 2020 - 19:32 BarryRothberg  
mods.36.StoneO.mmsmms mods.36.StoneO.mms manage 10 K 18 Aug 2020 - 19:32 BarryRothberg  
Topic revision: r65 - 05 Sep 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback